OpenStudy (anonymous):

Solve this system of equation 4x^2 − 9y^2 − 36 = 0 4x^2 − 36y = 0

OpenStudy (ajprincess):

subtract the first equation from the second equation. if you do so you will get a quadratic equation in terms of y. Then solve it to find y. Then using the value of y find x.

OpenStudy (anonymous):

4x^2 − 9y^2 − 36 = 0.......1 4x^2 − 36y = 0..........2 multiply 1 by -4 and add both 1 and 2 -12x^2+144=0

OpenStudy (anonymous):

find x, plug it into any of the 2 equations, then find y

OpenStudy (apoorvk):

Subtract equation 2 from equation 1 you get: \[-9y^2+36y-36=0\] now solve this by factorising, and get the two values of y. substitute and get the two values of x. and then do check if both sets of values satisfy both equations. post you answer to verify if you like!

OpenStudy (anonymous):

does y equal 2 and x equal square root of 18? or is x equal to + or - square root of 18?

OpenStudy (anonymous):

From apoorvk's work above...\[-9y^2+36y-36=0 \implies y^2-4y+4=0=(y-2)^2\implies y=2\]\[\implies x=\pm \sqrt{18}\implies (x,y)=(-\sqrt{18},2)~or~(\sqrt{18},2)\]

OpenStudy (ajprincess):

ya y=2 and x=+/-sqrt of 18.